OpenStudy (anonymous):

Consider the quadratic f(x)=x^2-2x-3 Find the x-intercepts of the graph of f(x). I just want to know the formula on how to find the x-intercepts.

OpenStudy (anonymous):

To find the x-intercept, you have to plug in y = 0 But it will be easier to find the x-intercept if we factor x^2-2x-3

OpenStudy (anonymous):

The factors would be (x-3)(x+1) right?

OpenStudy (anonymous):

@MeganXOXO

OpenStudy (anonymous):

Correct! And here's a small property: ab = 0, if a = 0, or b = 0, or if both a and b equal to 0 So we have (x-3)(x+1) = 0 x - 3 = 0 x + 1 = 0 solve for x. :)

OpenStudy (anonymous):

So the x-intercepts are (3,0) and (-1,0)? @MeganXOXO

OpenStudy (anonymous):

Correct! :)

OpenStudy (anonymous):

Thank you so much! :) Also to find the y-intercept is -3 because in standard form the c stands for the y-intercept, but I don't know how to figure that out by equation.. Do you know how? @MeganXOXO

OpenStudy (anonymous):

Yes, I do! @vpopstar13

OpenStudy (anonymous):

To find the y-intercept, you plug in x = 0 and solve for y So we had y = (x-3)(x+1) Plug in x = 0 y = (0-3)(0+1) y = (-3)(1) y = -3 (0,-3) Do you understand? @vpopstar13

OpenStudy (anonymous):

Ohhh, yes! Thank you so much again. @MeganXOXO

OpenStudy (anonymous):

I'm glad that I could help @vpopstar13 !

OpenStudy (anonymous):

I also have one more question... @MeganXOXO

OpenStudy (anonymous):

Sure @vpopstar13

OpenStudy (anonymous):

The vertex is the y-intercept right? But the problem is saying find the coordinates of the vertex of the graph.. I don't understand what it's saying. @MeganXOXO

OpenStudy (anonymous):

Nope @vpopstar13

OpenStudy (anonymous):

The vertex is -b/2a where the equation is ax^2 + bx + c =0

OpenStudy (anonymous):

So the vertex is the b in standard form? @MeganXOXO

OpenStudy (anonymous):

Wait I am so confused..

OpenStudy (anonymous):

Okay so the formula is a(x-h)^2 + k I just need to convert my standard form into vertex form, right? @MeganXOXO

OpenStudy (anonymous):

You can do that too! :)